### v1

# Computation of stresses in jammed packings modeled with Tresca friction

Abstract
This paper is interested in the computation of stresses within jammed packings of rigid polygonal cells. The cells are considered to follow a Tresca friction law. First, a constrained minimization problem is introduced where the friction energy is minimized while enforcing the non-interpenetration of neighboring cells as inequality constraint. The corresponding dual maximization problem is then deduced and its solution provides normal stresses at the interface between cells. Finally, lowest order Raviart–Thomas finite elements are used to reconstruct a consistent stress field by solving local problems. Numerical results are presented to showcase the consistency and robustness of the proposed methodology.
1 Introduction
Jammed particulate systems, such as granular materials, or masonry structures, exhibit complex mechanical behaviors arising from a disordered microstructure and a network of inter-particle contacts. When subjected to external loads, these systems support forces through a self-organized arrangement of contact stresses that ensure mechanical equilibrium. Understanding and accurately computing the stress fields within such jammed configurations is essential for predicting macroscopic mechanical response and failure mechanisms. Typical methods to compute stresses within granular material include Discrete Element Methods (DEM) as in [5, 12].
Granular materials inherently contain a lot of voids, which is an extra difficulty on its own. This paper is concerned with jammed packings that do not have such voids e.g. brick walls or vaults, see Figure 1.
(a) Brick wall
(b) Vault with a keystone
Figure 1: Examples of jammed packings
Let us denote by $\Omega\subset\mathbb{R}^{d}$, where $d=2,3$ a domain of interest (the jammed packing). We will consider that a packing is defined by a Voronoi mesh, denoted by $\mathcal{M}$, which perfectly matches the domain $\Omega$, see Figure 2.
Figure 2: Voronoi mesh $\mathcal{M}$ of $\Omega$.
Throughout this paper, we will refer to each element of the packing as a cell (of the mesh $\mathcal{M}$). In Figure 2, the blue dots represent the Voronoi centers of each cell. The interface between cells will then be the internal edges of the mesh $\mathcal{M}$.
For certain engineering applications (civil engineering in the case of Figure 1), the range of external loads that jammed packings are subjected to does not lead to a significant elastic deformation of the cells. Therefore, one can consider that the cells are fully rigid (i.e. they do not deform elastically). When the packing is submitted to external loads, even though it might not deform elastically, internal stresses are still present within $\Omega$. Under certain conditions, masonry structures [14] fall into that category. The stresses within such jammed packings can then be determined exclusively by friction forces. As [16] puts it, masonry structures are not accurately modeled by elasticity. We then decide to model the physics of jammed packings by the rigid friction of neighboring cells with a Tresca friction criterion. Plasticity in the joint and other physical phenomena [14] should be taken into account to produce a more physically realistic model. This work proposes a simplified model that allows to deduce stresses in jammed packings only from friction, without elastic effects.
Since [5], many DEM have computed friction forces by penalizing interpenetration. The friction forces are then deduced linearly from the interpenetration. Plasticity can also be taken into account in [5]. More recently, [13] has proposed an approach, using DEM, that does not allow interpenetration of neighboring cells and does not use elastic parameters. Within appropriate loading regimes, we believe this approach better captures the underlying physics of jammed packings. The main difference with our work is that [13] models large displacements whereas we are interested in the stresses of jammed packings.
This paper will first give a short reminder on Tresca friction. It will then present the reconstructions used to obtain a primal energy based on the Tresca criterion. The existence of minimizers of the primal energy is shown under appropriate hypotheses. Assuming the jammed packing is stable, the displacement of each cell should be zero (up to machine error). Thus, only solving the primal problem is not useful for the task at hand. The dual program is then introduced and classical results relating the primal and dual problem are recalled. The output of solving the dual problem is a collection of normal stresses at the edges of the mesh $\mathcal{M}$. Using this data, a consistent reconstruction of the stresses, is then presented. The reconstruction takes advantage of the fact that the lowest-order Raviart–Thomas elements $\mathbb{RT}_{0}$ are $H(\mathrm{div})$ conforming and that the associated degrees of freedom (dofs) are the normal stresses at the barycenter of the edges of the mesh. The idea is then to reconstruct in each cell $c\in\mathcal{M}$ a stress $\sigma_{c}\in\mathbb{R}^{d\times d}_{\text{sym}}$, the space of symmetric square tensors of size $d$, such that $\mathrm{div}(\sigma_{c})=0$ (in the absence of volume loads). Finally, numerical results are presented to show the consistency of the proposed methodology as well as its efficiency in computing the stresses for domains with numerous cells.
For simplicity, we consider in the rest of this paper that $d=2$. Extending the following methodology to $d=3$ is straightforward.
1.1 Tresca Friction
Let us illustrate the Tresca friction law in a very simple case. One can refer to the reference [3] for further detail. Let $\Omega\subset\mathbb{R}^{2}$ be an open bounded domain. We assume that the boundary of $\Omega$, written $\partial\Omega$, is Lipschitz and write as $\Gamma_{c}\subset\partial\Omega$ the contact surface. We write as $\bm{n}$ the outward unit normal to $\partial\Omega$ and as $\bm{t}=R(\frac{\pi}{2})\bm{n}$ the unit tangent to $\partial\Omega$, where $R(\frac{\pi}{2})$ is the 2d rotation matrix by an angle $\frac{\pi}{2}$. We write as $\bm{u}:\Omega\to\mathbb{R}^{2}$ the displacement field and $\sigma:\Omega\to\mathbb{R}^{2\times 2}_{\text{sym}}$ the stress tensor. We consider that the material in $\Omega$ is potentially in contact at $\Gamma_{c}$ with a fully rigid material (it does not deform). See figure 3 for a sketch of the setup.
[…]
Figure 3: Sketch of the contact between a rigid domain and $\Omega$.
Let us decompose the displacement and normal stress on $\Gamma_{c}$ as $\mathbf{u}=u_{\mathbf{n}}\mathbf{n}+\mathbf{u}_{\mathbf{t}}$ and $\sigma\mathbf{n}=\sigma_{\mathbf{n}}\mathbf{n}+\bm{\sigma}_{\mathbf{t}}$. The kinematic contact conditions at $\Gamma_{c}$ are:
[…]
The first equation in (1) prevents the penetration of $\Omega$ into the rigid domain, whereas the second describes that contact forces can only be repulsive. The third equation describes that two situations are possible at any point on $\bm{x}\in\Gamma_{c}$:
  • if the domain $\Omega$ is detached from the rigid support $u_{\bm{n}}(\bm{x})<0$, then necessarily $\sigma_{\bm{n}}=0$ (there cannot be a contact force),
  • if there is a negative contact force $\sigma_{\bm{n}}<0$, then necessarily the elastic body must stick to the support $u_{\bm{n}}=0$.
Note that (1) are sometimes called KKT conditions in the context of constrained optimization and in that context the normal stress $\sigma_{\bm{n}}$ is the Lagrange multiplier of the contact constraint. The equations in (1) are common to all contact laws and not specific to the Tresca friction law. The Tresca friction law writes on $\Gamma_{c}$ as
$\left\{\begin{aligned} &|\bm{\sigma}_{\mathbf{t}}|\leq s_{T}&\text{if }\mathbf{u}_{\bm{t}}=\bm{0},\\ &\bm{\sigma}_{\bm{t}}=-s_{T}\frac{\bm{u}_{\bm{t}}}{|\bm{u}_{\bm{t}}|}&\text{otherwise},\end{aligned}\right$. (2)
where $s_{T}>0$ is the Tresca friction coefficient. The energy dissipation due to the Tresca friction writes as
$\int_{\Gamma_{c}}s_{T}|\bm{u}_{\bm{t}}|.$ (3)
2 Discrete Model
We consider an open bounded polyhedral domain $\Omega\subset\mathbb{R}^{2}$. For simplicity, the domain will be subjected to normal stress (or Neumann) boundary conditions but no volume loads. Include volume loads would be a straightforward extension.
2.1 Setup
Let $\mathcal{C}$ be a collection of Voronoi cells perfectly matching $\Omega$. Let us write as $\mathcal{E}$ the set of edges of the mesh. The edges are partitioned as $\mathcal{E}=\mathcal{E}^{\partial}\cup\mathcal{E}^{i}$, where $\mathcal{E}^{\partial}$ are the external edges and $\mathcal{E}^{i}$ are the internal edges. For a cell $c\in\mathcal{C}$, we write the position of its Voronoi center by $\bm{x}_{c}$, which also represents its initial position. Each cell $c\in\mathcal{C}$ is composed of a collection of edges written $\mathcal{E}_{c}$. The cells are considered to be rigid, and thus we take the current position $\bm{z}\in\mathbb{P}^{0}(\mathcal{C})^{2}=:V$ as piecewise constant in each cell. The initial position of the cells is also represented by a piecewise constant functions $\bm{x}\in V$. The displacement $\bm{u}_{c}$ of a cell $c\in\mathcal{C}$ is defined as $\bm{u}_{c}:=\bm{z}_{c}-\bm{x}_{c}$. The collection of cell displacements is then stored in $\bm{u}\in V$. Since the cells are considered as rigid bodies, a complete description of the kinematics of cell $c$ would require a translation of the centroid, $\bm{u}_{c}$ and a rotation $\theta_{c}$, about that point. But, since in the present context, we consider jammed packings, the cells cannot rotate and we therefore do not consider any rotational degree of freedom. As a consequence, only force balance is imposed on each cell, while moment balance is not enforced.
Let us consider two cells $c_{+}$ and $c_{-}$, such that $e\in\mathcal{E}_{c_{+}}\cap\mathcal{E}_{c_{-}}$, as in Figure 4.
$\bm{x}_{c_{+}}$$\bm{x}_{c_{-}}$$\bm{n}_{e}$$e$
Figure 4: Notation for edge quantities
For $e\in\mathcal{E}^{i}$, one defines $\bm{n}_{e}$ as the unit normal to $e$ pointing from $c_{-}$ towards $c_{+}$. The jump operator is defined for $\bm{u}$ at $e$ by
$\bm{u}]_{e}:=\bm{u}_{c_{+}}-\bm{u}_{c_{-}}.$
For $e\in\mathcal{E}^{b}$, one defines $\bm{n}_{e}$ as the outward unit normal to $\partial\Omega$. Finally, for $e\in\mathcal{E}$, we define the tangent vector to $\bm{n}_{e}$ as $\bm{t}_{e}:=R(\frac{\pi}{2})\bm{n}_{e}$, where $R(\frac{\pi}{2})$ is the 2d rotation matrix by an angle $\frac{\pi}{2}$.
2.2 Facet values
In order to write the primal energy that we shall minimize, we need facet values computed from cell values. For a boundary edge $e\in\mathcal{E}^{\partial}$, we write the facet reconstruction operator as $\bm{R}_{e}(\bm{u})$. Since $e\in\mathcal{E}^{\partial}$, there exists a unique $c\in\mathcal{M}$ such that $e\subset\partial c$. The idea of the reconstruction is to simply use the displacement of the cell $c$ that contains the boundary facet $e$. Since the packing is jammed, cells cannot spin. Also, we consider the cells to be fully rigid, which makes this reconstruction consistent. One thus has
$\bm{R}_{e}(\bm{u}):=\bm{u}_{c}.$ (4)
2.3 Primal problem
We consider that the cells interact through Tresca friction forces, and that there can be no inter-penetration between the cells. Let us first describe the energy $E_{e}$ between the two neighboring cells $c_{1}$ and $c_{2}$ of Figure 4. We write the energy of the Tresca friction law (compare to (3)) for $e$ as
$E_{e}(\bm{u}):=s_{T}|e||[\bm{u}]_{e}\cdot\bm{t}_{e}|,$ (5)
In order to forbid inter-penetration between cells, we add the following constraint:
$\forall e\in\mathcal{E}^{i},\quad[\bm{u}]_{e}\cdot\bm{n}_{e}\geq 0.$ (6)
We also add a confinement term in order to keep the granular material compressed. It materializes as a Neumann boundary condition $\sigma\bm{n}=\bm{g}$ on $\partial\Omega$. Assume that $\bm{g}\in L^{2}(\partial\Omega)^{2}$ and let $e\in\mathcal{E}^{\partial}$. We define $\bm{g}_{e}:=|e|^{-1}\int_{e}\bm{g}$ as the average of $\bm{g}$ over $e$. Therefore, the work of external loads is defined as
$E_{\text{ext}}(\bm{u}):=\sum_{e\in\mathcal{E}^{\partial}}|e|\bm{g}_{e}\cdot\bm{R}_{e}(\bm{u}).$ (7)
The primal energy $E$ is then defined as
$E(\bm{u}):=\sum_{e\in\mathcal{E}^{i}}E_{e}(\bm{u})-E_{\text{ext}}(\bm{u}).$ (8)
Finally, since (8) has a translation invariance, we add the following equality constraint:
$\sum_{i=1}^{N}\bm{u}_{i}=0.$ (9)
We therefore define the primal problem as
$\min_{\bm{u}\in V}$ $E(\bm{u})$ (10)
subject to $\forall e\in\mathcal{E}^{i},\quad[\bm{u}]_{e}\cdot\bm{n}_{e}\geq 0,$
$\sum_{i=1}^{N}\bm{u}_{i}=0.$
2.4 Dual problem
The problem (10) is physically relevant but does not contain internal stresses. The normal stresses will appear as dual variables in the dual problem. The constrained problem (10) can be rewritten as an unconstrained minimization of a convex functional. Let us write as $I$ the indicator function defined as
$I(x)=\left\{\begin{aligned} &0&\text{ if }x\geq 0,\\ &+\infty&\text{ if }x<0.\end{aligned}\right$.
We can now define the proper convex function $\mathtt{E}:V\to\mathbb{R}\cup\{+\infty\}$ as
$\mathtt{E}(\bm{u})=E(\bm{u})+\sum_{e\in\mathcal{E}^{i}}I([\bm{u}]_{e}\cdot\bm{n}_{e}).$ (11)
Note that for simplicity, we neglect the equality constraint from (10) in the exposition of the dual problem.
Proposition 1.
Assume that $\bm{g}$ is such that $E(\bm{u})\to+\infty$ when $\|\bm{u}\|_{V}\to+\infty$. Then, the functional $\mathtt{E}$ defined in (11) admits a minimizer.
Remark 2.
Note that typical external loads $\bm{g}$ verifying the hypothesis of Proposition 1 contain compressive loads.
Proof of Proposition 1.
Let us define $\mathcal{V}:=\{\bm{v}\in V;\ \forall e\in\mathcal{E}^{i},\quad[\bm{u}]_{e}\cdot\bm{n}_{e}\geq 0\}$. $\mathcal{V}$ is a closed subspace of $V$ and thus is a vector space. Note that over $\mathcal{V}$, one has $\mathtt{E}=E$ and $\mathtt{E}$ is therefore continuous over $\mathcal{V}$. Using the hypothesis and the fact that $\mathcal{V}$ is finite-dimensional, once can deduce that $\mathtt{E}$ admits a global minimizer over $\mathcal{V}$. ∎
Let $W:=\mathbb{P}^{0}(\mathcal{E}^{i})$. We write dual variables as $\bm{f}\in W^{2}$. The dual energy $\mathtt{E}^{*}$ is then defined as the convex conjugate of $\mathtt{E}$ by
$\mathtt{E}^{*}(\bm{f})=\sup_{\bm{u}\in V}\left\{\langle\bm{f},\bm{u}\rangle-\mathtt{E}(\bm{u})\right\}.$ (12)
The dual problem is then defined as
$\max_{\bm{f}\in W^{2}}\mathtt{E}^{*}(\bm{f})$ (13)
Note that, because $\mathtt{E}$ has the physical dimension of an energy, then $\bm{f}$ is interpreted as a collection of forces associated to each internal edge, with the corresponding physical dimension. Let us recall the following classical result.
Proposition 3 (Weak Duality).
If $\bm{u}\in V$ is feasible for (10) and $\bm{f}$ is feasible for (13), then
$\mathtt{E}(\bm{u})\geq\mathtt{E}^{*}(\bm{f}).$
Note that Proposition 3 shows that, if Proposition 1 applies, then (13) admits a solution.
Remark 4.
Since $\mathtt{E}$ is not strictly convex, one cannot expect to have uniqueness of minimizers. A similar reasoning applies to $\mathtt{E}^{*}$ and (13).
Let $e\in\mathcal{E}^{i}$, we write $\bm{f}_{e}=(f^{n}_{e},f^{t}_{e})$. Assuming (13) has a feasible solution, the normal stress for a cell $c\in\mathcal{C}$ and an internal edge $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$ is thus defined as
$\bm{\lambda}_{e,c}:=|e|^{-1}(\bm{n}_{c}\cdot\bm{n}_{e})\left(f_{e}^{n}\bm{n}_{e}+f_{e}^{t}\bm{t}_{e}\right).$ (14)
Remark 5.
The principle of action-reaction is intrinsically taken into account here since for $c,c^{\prime}\in\mathcal{C}$ and $e\in\mathcal{E}_{c}\cap\mathcal{E}_{c^{\prime}}$, one has $\bm{\lambda}_{e,c}=-\bm{\lambda}_{e,c^{\prime}}$. That is due to the fact that one has, for instance, $\bm{n}_{c}=\bm{n}_{e}$ and $\bm{n}_{c^{\prime}}=-\bm{n}_{e}$.
3 Stress reconstruction
Solving (13) gives us, through (14), a normal stress in each internal edge $e\in\mathcal{E}^{i}$, which are completed by the normal stresses $\bm{g}$ imposed as Neumann boundary conditions on $\mathcal{E}^{\partial}$. Let us consider a cell $c\in\mathcal{C}$ as in Figure 5.
[…]
Figure 5: Stress reconstruction in an internal cell.
We define as $H(\mathrm{div};c)^{2}:=\{\sigma\in L^{2}(c)^{2\times 2};\mathrm{div}(\sigma)\in L^{2}(c)^{2}\}$. The associated norm is defined for $\sigma\in H(\mathrm{div};c)$ by $\|\sigma\|_{H(\mathrm{div};c)}^{2}:=\|\sigma\|_{L^{2}(c)}^{2}+\|\mathrm{div}(\sigma)\|_{L^{2}(c)}^{2}$. The idea of the reconstruction is to compute in each cell $c\in\mathcal{C}$, a field $\sigma_{c}\in H(\mathrm{div};c)^{2}$, so that $\mathrm{div}(\sigma_{c})=\bm{0}$ in $c$, since no volume load is applied to $\Omega$. This is performed by solving local problems with finite elements, as described in the following. Note that, if volumes loads were applied to $\Omega$, one could modify the continuity equation in order to take the volumes loads into account.
3.1 Geometric setup
Let $c\in\mathcal{C}$. We write as $\bm{n}_{c}$ the outward unit normal to $\partial c$ and the barycentre of the internal edge $e\in\mathcal{E}^{i}$ as $\bm{x}_{e}$. Since all cells are Voronoi, they are convex. As a consequence, its boundary $\partial c$ is Lipschitz [8]. Since any cell $c\in\mathcal{C}$ is convex, it can be triangulated by choosing an edge $e\in\mathcal{E}_{c}$ and the Voronoi center $\bm{x}_{c}$, as sketched in Figure 5. Let us write as $\mathcal{M}_{c}$ this triangulation of the cell $c$.
Let $\mathfrak{V}:=\mathbb{RT}_{0}(\mathcal{M}_{c})^{2}$, where $\mathbb{RT}_{0}(\mathcal{M}_{c})$ is the space of lowest-order Raviart–thomas elements, see [7]. We will use the Neumann boundary condition $\bm{g}$ and the solutions of (13) $\bm{\lambda}$ as Dirichlet boundary conditions. We define the solutions set as
$\mathfrak{V}_{D}:=\{\sigma\in\mathfrak{V};\,\forall e\in\mathcal{E}_{c}\cap\mathcal{E}^{\partial},\,\sigma(\bm{x}_{e})\bm{n}_{c}=\bm{g}_{e}\text{ and }\forall e\in\mathcal{E}_{c}\cap\mathcal{E}^{i},\,\sigma(\bm{x}_{e})\bm{n}_{c}=\bm{\lambda}_{e,c}\}.$
The associated homogeneous space is defined as
$\mathfrak{V}_{0}:=\{\sigma\in\mathfrak{V};\,\forall e\in\mathcal{E}_{c},\,\sigma(\bm{x}_{e})\bm{n}_{c}=\bm{0}\}.$
Note that, for $\sigma\in\mathfrak{V}$, one does not have $\sigma^{\mathsf{T}}=\sigma$. This condition will be imposed weakly in the following through a Lagrange multiplier. Let us define the space of Lagrange multipliers $Q:=\mathbb{P}^{0}(\mathcal{M}_{c})$, which is the space of piecewise constant functions.
3.2 Discrete problem
The formulation of this problem has been inspired by [2, 1]. For $\sigma,\tau\in H(\mathrm{div};c)^{2}$, we define the bilinear form
$a(\sigma,\tau):=\int_{c}\mathrm{div}(\sigma)\cdot\mathrm{div}(\tau).$
This bilinear form consists in a least-square formulation of the equation $\mathrm{div}(\sigma)=\bm{0}$. For $\sigma\in H(\mathrm{div};c)^{2}$ and $q\in L^{2}(c)$, we define the bilinear form
$b(\sigma,q):=\int_{c}\sigma:j(q),$
where $j(q):=\begin{pmatrix}0&q\\ -q&0\end{pmatrix}$ is a generic skew-symmetric matrix. Note that one has $|j(q)|=\sqrt{2}|q|$ when considering the Frobenius norm on matrices. This bilinear form is used to weakly impose $\sigma^{\mathsf{T}}=\sigma$ in $c$. The reconstruction consists in searching for $(\sigma,p)\in\mathfrak{V}_{D}\times Q$ such that
$a(\sigma,\tau)+b(\tau,p)=0,$ $\quad\forall\tau\in\mathfrak{V}_{0},$ (15)
$b(\sigma,q)=0,$ $\quad\forall q\in Q.$
Proposition 6.
There exists a unique solution of (15).
Remark 7.
Note that Proposition 6 shows that from a given $\bm{\lambda}\in W^{2}$, solution of (13), one can compute a unique $\sigma\in\mathfrak{V}_{D}$. But since $\bm{\lambda}\in W^{2}$, solution of (13) is not unique, the reconstructed stress $\sigma$ is not unique either.
Proof of Proposition 6.
In order to prove that (15) admits a unique solution, we want to use the BNB theorem for mixed systems, see [6, Theorem 2.34]. Let us first lift the Dirichlet boundary conditions. Using the surjectivity of the trace operator, see [7, Theorem 4.15], there exists $\mu\in H(\mathrm{div};c)^{2}$, such that for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{\partial}$, $\int_{e}\mu\bm{n}=|e|\bm{g}_{e}$ and for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$, $\int_{e}\mu\bm{n}=|e|\bm{\lambda}_{e,c}$. For $\tau\in\mathfrak{V}_{0}$, let us define the linear form
$F(\tau):=-\int_{c}\mathrm{div}(\mu)\cdot\mathrm{div}(\tau),$
which is continuous over $\mathfrak{V}_{0}$ and for $q\in Q$,
$G(q):=-\int_{c}\mu:j(q),$
which is linear continuous over $Q$. We now search for $(\sigma_{0},p)\in\mathfrak{V}_{0}\times Q$,
$a(\sigma_{0},\tau)+b(\tau,p)=L(\tau),$ $\quad\forall\tau\in\mathfrak{V}_{0},$
$b(\sigma_{0},q)=G(q),$ $\quad\forall q\in Q.$
Let us define $H_{0}(\mathrm{div};c)^{2}:=\{\sigma\in H(\mathrm{div};c)^{2};\sigma\bm{n}=\bm{0}\text{ on }\partial\Omega\}$. We propose to show first that $a$ is coervice over $Z:=H_{0}(\mathrm{div};c)^{2}\cap\{\sigma\in L^{2}(\Omega)^{2\times 2};\,\mathrm{curl}(\sigma)=\bm{0}\}$, where the 2d $\mathrm{curl}$ operator is applied row-wise. To prove that coercivity, we need to show that there exists $C>0$, such that for any $\sigma\in Z$, one has
$\|\sigma\|_{L^{2}(c)}\leq C\|\mathrm{div}(\sigma)\|_{L^{2}(c)}.$
Let us reason by density. Let $\varphi\in C^{\infty}_{c}(c)^{2\times 2}$, such that $\mathrm{curl}(\varphi)=\bm{0}$. Since $\varphi\in H^{1}_{0}(c)^{2\times 2}$, we can apply the Poincaré inequality and there exits $C>0$,
$\|\varphi\|_{L^{2}(c)}\leq C\|\nabla\varphi\|_{L^{2}(c)}.$
Using [4, Theorem 2.3] and the fact that $\mathrm{curl}(\varphi)=\bm{0}$, one has
$\|\varphi\|_{L^{2}(c)}\leq C\|\mathrm{div}(\varphi)\|_{L^{2}(c)}.$
We now consider a sequence $(\varphi_{n})_{n}\in C^{\infty}_{c}(c)^{2\times 2}$, such that $\mathrm{curl}(\varphi_{n})=\bm{0}$ and $\sigma\in Z$ such that $\|\varphi_{n}-\sigma\|_{H(\mathrm{div};c)}\to 0$, as $n\to+\infty$. Thus, as $n\to+\infty$, $\|\varphi_{n}-\sigma\|_{L^{2}(c)}\to 0$ and $\|\mathrm{div}(\varphi_{n}-\sigma)\|_{L^{2}(c)}$. One has
$\|\mathrm{div}(\sigma)\|_{L^{2}(c)}$ $=\|\mathrm{div}(\varphi_{n})|_{L^{2}(c)}+\|\mathrm{div}(\varphi_{n}-\sigma)|_{L^{2}(c)},$
$\geq\frac{1}{C}\|\varphi_{n}\|_{L^{2}(c)}+\|\mathrm{div}(\varphi_{n}-\sigma)|_{L^{2}(c)}\mathop{\longrightarrow}_{n\to+\infty}\frac{1}{C}\|\sigma\|_{L^{2}(c)}.$
Therefore,
$\|\sigma\|_{H(\mathrm{div};c)}^{2}=\|\mathrm{div}(\sigma)\|^{2}_{L^{2}(c)}+\|\sigma\|^{2}_{L^{2}(c)}\leq\left(1+C^{2}\right)\|\mathrm{div}(\sigma)\|^{2}_{L^{2}(c)}.$
Thus, one has
$a(\sigma,\sigma)=\|\mathrm{div}(\sigma)\|^{2}_{L^{2}(c)}\geq\frac{1}{1+C^{2}}\|\sigma\|_{H(\mathrm{div};c)}^{2},$
which shows the coercivity over $Z$. To finish, we note that $\mathfrak{V}_{0}$ is conforming in $Z$ in the sense that $\mathfrak{V}_{0}\subset Z$.
In order to apply the BNB theorem for mixed formulation [6, Theorem 2.34], it remains to show the second condition of the theorem. Let $q\in Q$ and $\tau\in\mathfrak{V}_{0}$. One has
$b(\tau,q)=\sum_{T\in\mathcal{M}_{c}}\int_{T}\tau:j(q)$ $=\sum_{T\in\mathcal{M}_{c}}|T|\left<\tau\right>_{T}:j(q),$
$\leq\sum_{T\in\mathcal{M}_{c}}|T|\left|\left<\tau\right>_{T}\right|\sqrt{2}|q|,$
$\leq\sqrt{2}\left(\sum_{T\in\mathcal{M}_{c}}|T|\left|\left<\tau\right>_{T}\right|^{2}\right)^{\frac{1}{2}}\left(\sum_{T\in\mathcal{M}_{c}}|T||q|^{2}\right)^{\frac{1}{2}},$
$=\sqrt{2}\|\left<\tau\right>_{T}\chi_{T}\|_{L^{2}(c)}\|q\|_{L^{2}(c)}$
where $\left<\tau\right>_{T}$ is the average over $T$ and $\chi_{T}$ is the indicator function over $T$. Considering $\tau\in\mathfrak{V}_{0}\cap\mathbb{P}^{0}(\mathcal{M}_{c})^{2\times 2}$, one has $\|\tau\|_{H(\mathrm{div};c)}=\|\tau\|_{L^{2}(c)}$, and thus
$\frac{b(\tau,q)}{\|\tau\|_{H(\mathrm{div};c)}}=\sqrt{2}\|q\|_{L^{2}(c)}.$
The supremum over $\tau\in\mathfrak{V}_{0}$ is thus achieved and one has
$\inf_{q\in Q}\sup_{\tau\in\mathfrak{V}_{0}}\frac{b(\tau,q)}{\|\tau\|_{H(\mathrm{div};c)}\|q\|_{L^{2}(c)}}\geq\sqrt{2},$
which finishes the proof. ∎
4 Numerical results
We first describe how the normal stresses are computed in practice. Then, we present a few tests to ensure the consistency of the proposed method. Finally, we show that the method can handle any Voronoi meshes.
4.1 Linear program
Let us rewrite (10) as a linear program. This allows us to use very efficient linear solvers. For each internal edge $e\in\mathcal{E}^{i}$, we add a new variable $v_{e}$ such that $v_{e}\geq|[\bm{u}]_{e}\cdot\bm{t}_{e}|$. Note that this inequality constraint can be written with two linear constraints. Let us write $v:=(v_{e})_{e}\in W$. Therefore, the linear program is now
$\min_{(\bm{u},v)\in V\times W}$ $\sum_{e\in\mathcal{E}^{i}}s_{T}|e|v_{e}-E_{\text{ext}}(\bm{u})$ (16)
subject to $\forall e\in\mathcal{E}^{i},\quad[\bm{u}]_{e}\cdot\bm{n}_{e}\geq 0,$
$\forall e\in\mathcal{E}^{i},\quad v_{e}\geq[\bm{u}]_{e}\cdot\bm{t}_{e},$
$\forall e\in\mathcal{E}^{i},\quad v_{e}\geq-[\bm{u}]_{e}\cdot\bm{t}_{e},$
$\sum_{i=1}^{N}\bm{u}_{i}=0.$
4.2 Implementation
The implementation is performed in Python. The package pyvoro, which uses the library Voro++ [15] is used to generate the Voronoi mesh. The package cvxopt [17] is used to solve (16) and provide the normal stresses by computing the dual linear program. Finally, Firedrake [9] is used to solve (15). The implementation is available in [11].
4.3 Verification tests
For the first test case, we consider $\Omega=(0,1)^{2}$ and the mesh sketched in Figure 6(a).
(a) Mesh $\mathcal{M}$ of $\Omega$
(b) Reconstructed stresses $\sigma_{ij}$.
Figure 6: Verification test 1
The Tresca friction coefficient is taken as $s_{T}:=10$. Let $S=-\begin{pmatrix}1&1\\ 1&1\\ \end{pmatrix}$. The boundary condition applied to the domain is $\bm{g}:=S\bm{n}$. Figure 6(b) shows the resconstructed stress $\sigma$. Note that $\sigma_{11}=\sigma_{22}=\sigma_{12}=\sigma_{21}$ in Figure 6(b) and therefore only one picture is shown. The result we obtain is consistent with the imposed boundary condition since we expect $\sigma=S$ for this test case.
For the second test case, we consider the same domain, material parameters and boundary conditions. However, we consider the mesh sketched in Figure 7(a).
(a) Mesh $\mathcal{M}$ of $\Omega$
(b) Reconstructed stress $\sigma_{ij}$.
Figure 7: Verification test 2
Figure 7(b) shows the reconstructed stress $\sigma$. The result we obtain is consistent with the imposed boundary condition since we expect $\sigma=S$ for this test case too.
4.4 Shear test
We consider $\Omega=(0,1)^{2}$, the mesh shown in Figure 8(a) and the boundary conditions sketched in Figure 8(b).
(a) Mesh $\mathcal{M}$ of $\Omega$.
$\Omega$$\bm{g}_{e}=\bm{e}_{1}$$\bm{g}_{e}=-\bm{e}_{1}$
(b) Boundary conditions
Figure 8: Shear test
We write the usual direct orthonormal basis of $\mathbb{R}^{2}$ as $(\bm{e}_{1},\bm{e}_{2})$. This test consists in shear loading a brick wall. Note that the weight of the bricks is ignored here. The computed stresses are presented in Figure 9.
(a) Reconstructed stress $\sigma_{11}$
(b) Reconstructed stress $\sigma_{12}$
(c) Reconstructed stress $\sigma_{21}$
(d) Reconstructed stress $\sigma_{22}$
Figure 9: Shear test
Note that the stresses are discontinuous at the interfaces between bricks since one only has $\sigma\in H(\mathrm{div};c)$, for all $c\in\mathcal{M}$. Also, one notices that the stress tensors are not completely symmetric.
4.5 Homogeneous compression test
We consider $\Omega=(0,1)^{2}$ and the mesh sketched in Figure 10.
Figure 10: Homogeneous compression test: mesh $\mathcal{M}$ of $\Omega$.
It contains 60 Voronoi cells. The boundary conditions consists in imposing $\bm{g}:=-\bm{n}$ on $\partial\Omega$. The results are shown in Figure 11.
(a) Reconstructed stress $\sigma_{11}$
(b) Reconstructed stress $\sigma_{12}$
(c) Reconstructed stress $\sigma_{21}$
(d) Reconstructed stress $\sigma_{22}$
Figure 11: Homogeneous compression test
We notice that the recomputed stress $\sigma$ in Figure 11 is not homogeneous. Let us consider the constant stress $\sigma=-I_{2}$, since it would seem as a good guess for this test case. Upon recomputing the forces $\bm{f}$ from (14) and $\sigma=-I_{2}$, we notice that $\sigma=-I_{2}$ is also an admissible maximizer of (13) for this test case. It is just not the solution found by the algorithm used in cvxopt. Indeed, this configuration is statically indeterminate and there are several force configurations satisfying the conditions for equilibrium. Note that some stress concentration is observed in Figure 11 since $|\sigma_{ij}|$ can get larger than $|\bm{g}|=1$. Finally, even though the boundary condition is a homogeneous compression, we observe that several cells are in tension. We believe that even though this result sounds counter-intuitive, it is logical since the friction between neighboring cells can lead to having cells under local tension. This is a well documented phenomena. Local tension is required for a crack to open which itself is required for a crack to propagate. In [10], local tension is observed in rocks under compression when they fracture.
Data Availability Statement